OpenStudy (anonymous):

The volume of a rectangular prism is 300 ft3. If the area of the base is 12 ft2, what is its height?

OpenStudy (johnweldon1993):

The height of a rectangular prism...can be evaluated by dividing the volume by the area of the base

OpenStudy (anonymous):

25

OpenStudy (anonymous):

V=LBH BUT V= 300, LB=12 THEREFORE 300=12H H=300/12 =25

OpenStudy (johnweldon1993):

yes @Brotherman but also remember the units...ft^3/ft^2 = ft...so it would be 25ft makes sense too...because area would be Length * width.....and volume is length * width * height.... so the volume = area * height = 12 ft^2 * 25ft = 300 ft^3 gotta make sure the units are correct to check

OpenStudy (anonymous):

yeah, the ft3/ft2 will give us ft